my question is ;-
explain graphically
3x+15y=24
3x-8y =24
plot points on graph and draw a straight line passing through that two points

Answer by ichudov(507)

you need to convert these equations to form y=...
3x+15y=24 converts to y=-(3/15)*x+24/15
3x-8y =24 converts to y=(3/8)*x-24/8
You can now plot these curves by plotting opints and drawing a straight line: